# Contraction of Private Quantum Channels and Private Quantum Hypothesis Testing

### v1

@@ Abstract
A quantum generalized divergence by definition satisfies the data-processing inequality; as such, the relative decrease in such a divergence under the action of a quantum channel is at most one. This relative decrease is formally known as the contraction coefficient of the channel and the divergence. Interestingly, there exist combinations of channels and divergences for which the contraction coefficient is strictly less than one. Furthermore, understanding the contraction coefficient is fundamental for the study of statistical tasks under privacy constraints. To this end, here we establish upper bounds on contraction coefficients for the hockey-stick divergence under privacy constraints, where privacy is quantified with respect to the quantum local differential privacy (QLDP) framework, and we fully characterize the contraction coefficient for the trace distance under privacy constraints. With the machinery developed, we also determine an upper bound on the contraction of both the Bures distance and quantum relative entropy relative to the normalized trace distance, under QLDP constraints. Next, we apply our findings to establish bounds on the sample complexity of quantum hypothesis testing under privacy constraints. Furthermore, we study various scenarios in which the sample complexity bounds are tight, while providing order-optimal quantum channels that achieve those bounds. Lastly, we show how private quantum channels provide fairness and Holevo information stability in quantum learning settings.
@@ I Introduction
With the development of quantum technologies and generation of quantum data, ensuring privacy of quantum systems is an important research question. To this end, statistical privacy frameworks for quantum data have been developed recently, which are generalizations of classical statistical privacy frameworks [11, 12, 13, 14]. Quantum local differential privacy (QLDP), which is a generalization of LDP, ensures that two distinct quantum states passed through a quantum channel are hard to distinguish by a measurement [13]. Several works have explored applications of these frameworks in quantum machine learning [15, 16, 17, 18, 19, 20, 21].
@@
Symmetric hypothesis testing under LDP constraints (when one has access to privatized data samples in contrast to the data from original distributions) has been extensively studied in a chain of works [8, 9, 33, 10, 34]. In [8], a contraction bound of the Kullback–Leibler divergence relative to the total-variation distance has been derived, which implies that one needs
@@ I-A Contributions
Together with the tools developed, we provide upper and lower bounds on the sample complexity when symmetric hypothesis testing is carried out with access to privatized quantum states (Theorem 4). The upper and lower bounds in Theorem 4 do not match in general: the upper bound scales as $1/[T(\rho,\sigma)]^{2}$ whereas the lower bound scales as $1/T(\rho,\sigma)$. For the specific case in which the states are orthogonal, we prove that the sample complexity scales as $\Theta\!\left(\left[(e^{\varepsilon}+1)/(e^{\varepsilon}-1)\right]^{2}\right)$, showcasing the cost of privacy (Corollary 2). Furthermore, we consider a specific setting where the private mechanism is chosen from a special class of channels (defined in (159)) depending on the states used in the hypothesis testing task. There, we establish that the sample complexity is characterized by $\Theta\!\left(\left({1}/\left[{\varepsilon T(\rho,\sigma)}\right]\right)^{2}\right)$ for $\varepsilon<1$. In this case, we show that a channel belonging to QLDP mechanisms proposed in Proposition 1 achieves order optimality. We also find that under certain low privacy regimes (higher $\varepsilon$) the private sample complexity is behaving similar to then non-private setting (Proposition 6). We extend our analysis to private sample complexity bounds of multiple hypothesis testing and asymmetric hypothesis testing in Remark 8 and Remark 9, respectively.
Moreover, we also precisely characterize the privatized contraction coefficient of the trace distance under $(\varepsilon,\delta)$-QLDP privacy constraints in Theorem 5, via a different proof argument for converse when compared to the proof of Theorem 2. To this end, we show that the privatized contraction coefficient under these privacy constraints evaluates to $(e^{\varepsilon}-1+2\delta)/(e^{\varepsilon}+1)$, generalizing the finding in Theorem 2.
@@ I-B Organization
The rest of our paper is organized as follows. First we introduce notation and required background in Section II. We study privatized contraction coefficients under local-quantum privacy constraints in Section III, which may be of independent interest for a wide spectrum of statistical tasks under privacy constraints. Using the tools developed in Section III, we present how the sample complexity of quantum symmetric hypothesis testing scales with the privacy parameters and the states of interest (see Section IV). In addition, we also characterize the privatized contraction coefficient of trace distance under $(\varepsilon,\delta)$-QLDP privacy constraints in Section V, generalizing some results presented in Section III. Furthermore, we explore applications in ensuring fairness and stability in learning algorithms in Section VI. Lastly, we conclude our paper in Section VII with a summary and some future directions to explore.
@@ II-A Notation
A quantum state $\rho_{R}\in\mathcal{L}(\mathcal{H}_{R})$ on $R$ is a PSD, unit-trace operator acting on $\mathcal{H}_{R}$. We denote the set of all quantum states in $\mathcal{L}(\mathcal{H}_{R})$ by $\mathcal{D}(\mathcal{H}_{R})$. A state $\rho_{R}$ of rank one is called pure, and we may choose a normalized vector $|\psi\rangle\in\mathcal{H}_{R}$ satisfying $\rho_{R}=|\psi\rangle\!\langle\psi|$ in this case. Otherwise, $\rho_{R}$ is called a mixed state. A quantum channel $\mathcal{N}\colon\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is a linear, completely positive and trace-preserving (CPTP) map from $\mathcal{L}(\mathcal{H}_{A})$ to $\mathcal{L}(\mathcal{H}_{B})$. We often use the shorthand $\mathcal{N}_{A\to B}$ for such a channel, and we use the notation CPTP to denote the set of all quantum channels. We denote the adjoint of $\mathcal{N}$ by $\mathcal{N}^{\dagger}$. A measurement of a quantum system $R$ is described by a positive operator-valued measure (POVM) $\{M_{y}\}_{y\in\mathcal{Y}}$, which is defined to be a collection of PSD operators satisfying $\sum_{y\in\mathcal{Y}}M_{y}=I_{R}$, where $\mathcal{Y}$ is a finite alphabet. According to the Born rule, after applying the above POVM to $\rho\in\mathcal{D}(\mathcal{H}_{R})$, the probability of observing the outcome $y$ is given by $\mathrm{Tr}\!\left[M_{y}\rho\right]$.
@@ II-B Quantum Divergences
It generalizes the total-variation distance between two probability distributions. The quantum hockey-stick divergence is defined as [38]
@@
$D(\rho\|\sigma)\coloneqq\begin{cases}\mathrm{Tr}\!\left[\rho(\log\rho-\log\sigma)\right]&\text{ if }\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)\\ +\infty&\text{ otherwise }\end{cases},$ (8)
@@ II-C Privacy Models
concluding the claim. For the case where $\delta=0$, $\varepsilon$-QLDP is equivalent to
@@
Remark 2 (Connection to Max-Relative Entropy).
@@ III Contraction under Quantum Privacy Constraints
where $p\in[0,1]$ and $d$ is the dimension of the Hilbert space on which $\rho$ acts. In [14, Remark 9], a mechanism achieving $\varepsilon$-QLDP based on a depolarizing channel was presented. However, the parameters therein depend on $d$. Here, we present a mechanism that employs a measurement channel along with the depolarizing channel to achieve $\varepsilon$-QLDP, where the parameters are independent of $d$.
@@ III-A Contraction Coefficients under Privacy Constraints
Here we define privatized contraction coefficients as follows:
@@
Next we provide an upper bound on the privatized contraction coefficient defined in (38) for $\gamma\in[1,e^{\varepsilon}]$. Note that $\eta_{E_{\gamma}}^{\varepsilon}=0$ for $\gamma\geq e^{\varepsilon}$, as a consequence of (17) and the fact that $E_{\gamma}$ is monotone non-increasing in $\gamma$ [42, Lemma 4.2]. The privatized contraction coefficient of the hockey-stick divergence in the classical privacy setting (i.e., analogous to Theorem 1) was established in [43, Theorem 1], and our proof below is inspired by their proof.
@@
The fact that (50) and (52) are equal was stated in the proof of [43, Theorem 1]. Here we provide a short proof for convenience: The first equality above follows from the fact that the states $\omega_{p}$ and $\omega_{q}$ correspond to Bernoulli distributions with parameters $p$ and $q$, respectively, and the reasoning given below: $E_{e^{\varepsilon}}\!\left(\omega_{p}\|\omega_{q}\right)=0$ is equivalent to $p\leq e^{\varepsilon}q$ and $(1-p)\leq e^{\varepsilon}(1-q)$, and $E_{e^{\varepsilon}}\!\left(\omega_{q}\|\omega_{p}\right)=0$ is equivalent to $q\leq e^{\varepsilon}p$ and $(1-q)\leq e^{\varepsilon}(1-p)$. Without loss of generality, suppose that $q\leq p$. Then all of these inequalities can be written in the following compact form: $q\leq p\leq\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}$. The last equality follows by solving the linear program. Indeed, we can eliminate $p$ by noting that $\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}\leq 1$ and thus we can maximize its value by picking $p=\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}$. From there, observe that $qe^{\varepsilon}\leq\ qe^{-\varepsilon}+1-e^{-\varepsilon}$ is equivalent to $q\leq 1/(e^{\varepsilon}+1)$, and $qe^{\varepsilon}\geq\ qe^{-\varepsilon}+1-e^{-\varepsilon}$ is equivalent to $q\geq 1/(e^{\varepsilon}+1)$. With this, $p-\gamma q$ is supremized at $q=1/(e^{\varepsilon}+1)$. This concludes the proof. ∎
Hockey-stick divergence is defined as (7) for $\gamma\geq 1$. Now, similar to the classical definition of hockey-stick divergence (see [43, Eq.12]), we extend the definition of hockey-stick divergence for $\gamma\geq 0$ as follows:
@@
where $\rho$ and $\sigma$ are states and $(x)_{+}\coloneqq\max\{0,x\}$. This definition recovers (7) for $\gamma\geq 1$ since $(1-\gamma)_{+}=0$.
@@
Furthermore, using the monotoncity of $\gamma^{\prime}\mapsto E_{\gamma^{\prime}}$ for $\gamma^{\prime}\geq 1$, for $\gamma\leq e^{-\varepsilon}$
@@
Next, we quantify the contraction of hockey-stick divergence for $\gamma\in[e^{-\varepsilon},1]$ similar to Theorem 1 for $\gamma\in[1,e^{\varepsilon}]$.
@@
Let $\gamma\in[e^{-\varepsilon},1]$, $\mathcal{A}\in\mathcal{B}^{\varepsilon}$, and $\rho$ and $\sigma$ be states. Then, we have that
@@
where $E_{\gamma}(\cdot\|\cdot)$ is defined in (53). Furthermore, for $\gamma\in[e^{-\varepsilon},1]$ privatized contraction coefficient of hockey-stick divergence $\eta^{\varepsilon}_{E_{\gamma}}$ satisfies the following inequality:
@@
Since $\gamma\in[e^{-\varepsilon},1]$, $1/\gamma\in[1,e^{\varepsilon}]$. By applying Lemma 1, we obtain the following:
@@
In fact, Theorem 1 and Corollary 1 strengthen the previously known upper bound on the privatized contraction coefficient for the normalized trace distance (i.e., $\eta^{\varepsilon}_{T}$ by choosing $\gamma=1$), as recalled in (37). Next, Theorem 2 states that the upper bound from Theorem 1, for trace distance, is indeed tight. Under $\varepsilon$-classical local differential privacy (recall Definition 1), Theorem 2 was presented as [7, Corollary 2.9]. Our method of proof is different from that presented for [7, Corollary 2.9].
@@
$D_{f}(\rho\|\sigma)$ $=\int_{0}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)\,\mathsf{d}\gamma+\int_{0}^{1}f^{\prime\prime}(\nu)E_{\nu}(\rho\|\sigma)\,\mathsf{d}\nu$ (104)

### v2

@@ Abstract
A quantum generalized divergence by definition satisfies the data-processing inequality; as such, the relative decrease in such a divergence under the action of a quantum channel is at most one. This relative decrease is formally known as the contraction coefficient of the channel and the divergence. Interestingly, there exist combinations of channels and divergences for which the contraction coefficient is strictly less than one. Furthermore, understanding the contraction coefficient is fundamental for the study of statistical tasks under privacy constraints. To this end, here we establish upper bounds on contraction coefficients for the hockey-stick divergence under privacy constraints, where privacy is quantified with respect to the quantum local differential privacy (QLDP) framework, and we fully characterize the contraction coefficient for the trace distance under privacy constraints. Using the machinery developed, we also determine an upper bound on the contraction of both the Bures distance and quantum relative entropy relative to the normalized trace distance, under QLDP constraints. Next, we apply our findings to establish bounds on the sample complexity of quantum hypothesis testing under privacy constraints. Furthermore, we study various scenarios in which the sample complexity bounds are tight, while providing order-optimal quantum channels that achieve those bounds. Lastly, we show how private quantum channels provide fairness and Holevo information stability in quantum learning settings.
Index Terms:
Contraction coefficients, cost of privacy, hypothesis testing, sample complexity, quantum divergences, quantum privacy
@@ I Introduction
With the development of quantum technologies and generation of quantum data, ensuring privacy of quantum systems is an important research direction. To this end, statistical privacy frameworks for quantum data have been developed recently, which are generalizations of classical statistical privacy frameworks [11, 12, 13, 14]. Quantum local differential privacy (QLDP), which is a generalization of LDP, ensures that two distinct quantum states passed through a quantum channel are hard to distinguish by a measurement [13]. Several works have explored applications of these frameworks in quantum machine learning [15, 16, 17, 18, 19, 20, 21].
@@
Symmetric hypothesis testing under LDP constraints (when one has access to privatized data samples in contrast to the data from the original distributions) has been extensively studied in a chain of works [8, 9, 33, 10, 34]. In [8], a contraction bound for the Kullback–Leibler divergence relative to the total-variation distance has been derived, which implies that one needs
@@ I-A Contributions
Together with the tools developed, we provide upper and lower bounds on the sample complexity when symmetric hypothesis testing is carried out with access to privatized quantum states (Theorem 4). The upper and lower bounds in Theorem 4 do not match in general: the upper bound scales as $1/[T(\rho,\sigma)]^{2}$ whereas the lower bound scales as $1/T(\rho,\sigma)$. For the specific case in which the states are orthogonal, we prove that the sample complexity scales as $\Theta\!\left(\left[(e^{\varepsilon}+1)/(e^{\varepsilon}-1)\right]^{2}\right)$, showcasing the cost of privacy (Corollary 2). Furthermore, we consider a specific setting where the private mechanism is chosen from a special class of channels (defined in (163)) depending on the states used in the hypothesis testing task. There, we establish that the sample complexity is characterized by $\Theta\!\left(\left({1}/\left[{\varepsilon T(\rho,\sigma)}\right]\right)^{2}\right)$ for $\varepsilon<1$. In this case, we show that a channel belonging to QLDP mechanisms proposed in Proposition 1 achieves order optimality. We also find that under certain low-privacy regimes (higher $\varepsilon$) the private sample complexity behaves similarly to the non-private setting (Proposition 7). We extend our analysis to private sample complexity bounds for multiple hypothesis testing and asymmetric hypothesis testing in Remark 8 and Remark 9, respectively.
Moreover, we also precisely characterize the privatized contraction coefficient of the trace distance under $(\varepsilon,\delta)$-QLDP privacy constraints in Theorem 5, via a different proof argument for the converse, when compared to the proof of Theorem 2. To this end, we show that the privatized contraction coefficient under these privacy constraints evaluates to $(e^{\varepsilon}-1+2\delta)/(e^{\varepsilon}+1)$, generalizing the finding in Theorem 2.
@@
Note on related work: During the development of our paper, we became aware of independent and concurrent work, now available at [36], which also considers sample complexity of hypothesis testing under $\varepsilon$-local differential privacy constraints.
@@ I-B Organization
The rest of our paper is organized as follows. First we introduce notation and required background in Section II. We study privatized contraction coefficients under local-quantum privacy constraints in Section III, which may be of independent interest for a wide spectrum of statistical tasks under privacy constraints. Using the tools developed in Section III, we present how the sample complexity of quantum symmetric hypothesis testing scales with the privacy parameters and the states of interest (see Section IV). In addition, we also characterize the privatized contraction coefficient for trace distance under $(\varepsilon,\delta)$-QLDP privacy constraints in Section V, generalizing some results presented in Section III. Furthermore, we explore applications in ensuring fairness and stability in learning algorithms in Section VI. Lastly, we conclude our paper in Section VII with a summary and some future directions to explore.
@@ II-A Notation
A quantum state $\rho_{R}\in\mathcal{L}(\mathcal{H}_{R})$ on $R$ is a PSD, unit-trace operator acting on $\mathcal{H}_{R}$. We denote the set of all quantum states in $\mathcal{L}(\mathcal{H}_{R})$ by $\mathcal{D}(\mathcal{H}_{R})$. A state $\rho_{R}$ of rank one is called pure, and we may choose a normalized vector $|\psi\rangle\in\mathcal{H}_{R}$ satisfying $\rho_{R}=|\psi\rangle\!\langle\psi|$ in this case. Otherwise, $\rho_{R}$ is called a mixed state. A quantum channel $\mathcal{N}\colon\mathcal{L}(\mathcal{H}_{A})\to\mathcal{L}(\mathcal{H}_{B})$ is a linear, completely positive, and trace-preserving (CPTP) map from $\mathcal{L}(\mathcal{H}_{A})$ to $\mathcal{L}(\mathcal{H}_{B})$. We often use the shorthand $\mathcal{N}_{A\to B}$ for such a channel, and we use the notation CPTP to denote the set of all quantum channels. We denote the Hilbert–Schmidt adjoint of $\mathcal{N}$ by $\mathcal{N}^{\dagger}$. A measurement of a quantum system $R$ is described by a positive operator-valued measure (POVM) $\{M_{y}\}_{y\in\mathcal{Y}}$, which is defined to be a collection of PSD operators satisfying $\sum_{y\in\mathcal{Y}}M_{y}=I_{R}$, where $\mathcal{Y}$ is a finite alphabet. According to the Born rule, after applying the above POVM to $\rho\in\mathcal{D}(\mathcal{H}_{R})$, the probability of observing the outcome $y$ is given by $\mathrm{Tr}\!\left[M_{y}\rho\right]$.
@@ II-B Quantum Divergences
It generalizes the total-variation distance between two probability distributions. For $\gamma\geq 1$, the quantum hockey-stick divergence is defined as [39]
@@
$D(\rho\|\sigma)\coloneqq\begin{cases}\mathrm{Tr}\!\left[\rho(\ln\rho-\ln\sigma)\right]&\text{ if }\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)\\ +\infty&\text{ otherwise }\end{cases},$ (8)
@@ II-C Privacy Models
concluding the claim. For the case in which $\delta=0$, $\varepsilon$-QLDP is equivalent to
@@
Remark 2 (Connection to Max-Relative Entropy and Datta–Leditzky Divergence).
@@
By adapting [14, Proposition 1], $(\varepsilon,\delta)$-QLDP is also equivalent to the following condition:
$\sup_{\rho,\sigma\in\mathcal{D}(\mathcal{H})}\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)\leq\varepsilon,$ (22)
where
$\overline{\mathsf{D}}^{\delta}(\rho\|\sigma)\coloneqq\ln\inf\left\{\lambda\geq 0:\operatorname{Tr}[\left(\rho-\lambda\sigma\right)_{+}]\leq\varepsilon\right\}.$ (23)
$\sup_{\rho,\sigma\in\mathcal{D}(\mathcal{H})}\overline{\mathsf{D}}^{\delta}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\sup_{\varphi_{1}\perp\varphi_{2}}\overline{\mathsf{D}}^{\delta}(\mathcal{A}(\varphi_{1})\|\mathcal{A}(\varphi_{2}))\leq\varepsilon,$ (24)
where $\varphi_{1}$ and $\varphi_{2}$ are orthogonal pure states. The proof of this fact is in Section -B. For $\delta=0$, this can also be deduced from [43, Theorem 5.7].
For further introductions to statistical privacy frameworks including quantum differential privacy, and quantum pufferfish privacy and their potential applications, we refer to [13, 14].
@@ III Contraction under Quantum Privacy Constraints
where $p\in[0,1]$ and $d$ is the dimension of the Hilbert space on which $\rho$ acts. A depolarization mechanism achieving standard quantum differential privacy with a neighboring relation declared by the closeness of normalized trace distance was presented in [11, 13]. Considering the worst case scenario where the normalized trace distance constraint evaluates to one, a mechanism achieving $\varepsilon$-QLDP based on a depolarizing channel can be obtained, as shown in [14, Remark 9]. However, the parameters therein depend on $d$.
In [44, Lemma 5.2], a general recipe to derive privacy mechanisms by applying a POVM followed by a classical privacy mechanism, has been studied. The next proposition derives a QLDP mechanism using a similar technique, where a two-outcome POVM is followed by the application of a depolarizing mechanism. In this mechanism, the parameters are independent of the dimension $d$ of the input states.
@@ III-A Contraction Coefficients under Privacy Constraints
Here we define a privatized contraction coefficient for a generalized divergence $\boldsymbol{D}$ as follows:
@@
Next we provide an upper bound on the privatized contraction coefficient defined in (41) for $\gamma\in[1,e^{\varepsilon}]$. Note that $\eta_{E_{\gamma}}^{\varepsilon}=0$ for $\gamma\geq e^{\varepsilon}$, as a consequence of (17) and the fact that $E_{\gamma}$ is monotone non-increasing in $\gamma$ [45, Lemma 4.2]. The privatized contraction coefficient of the hockey-stick divergence in the classical privacy setting (i.e., analogous to Theorem 1) was established in [46, Theorem 1], and our proof below is inspired by their proof.
@@
The fact that (53) and (55) are equal was stated in the proof of [46, Theorem 1]. Here we provide a short proof for convenience: The first equality above follows from the fact that the states $\omega_{p}$ and $\omega_{q}$ correspond to Bernoulli distributions with parameters $p$ and $q$, respectively, and the reasoning given below: $E_{e^{\varepsilon}}\!\left(\omega_{p}\|\omega_{q}\right)=0$ is equivalent to $p\leq e^{\varepsilon}q$ and $(1-p)\leq e^{\varepsilon}(1-q)$, and $E_{e^{\varepsilon}}\!\left(\omega_{q}\|\omega_{p}\right)=0$ is equivalent to $q\leq e^{\varepsilon}p$ and $(1-q)\leq e^{\varepsilon}(1-p)$. Without loss of generality, suppose that $q\leq p$. Then all of these inequalities can be written in the following compact form: $q\leq p\leq\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}$. The last equality follows by solving the linear program. Indeed, we can eliminate $p$ by noting that $\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}\leq 1$ and thus we can maximize its value by picking $p=\min\{qe^{\varepsilon},\ qe^{-\varepsilon}+1-e^{-\varepsilon}\}$. From there, observe that $qe^{\varepsilon}\leq\ qe^{-\varepsilon}+1-e^{-\varepsilon}$ is equivalent to $q\leq 1/(e^{\varepsilon}+1)$, and $qe^{\varepsilon}\geq\ qe^{-\varepsilon}+1-e^{-\varepsilon}$ is equivalent to $q\geq 1/(e^{\varepsilon}+1)$. With this, $p-\gamma q$ is supremized at $q=1/(e^{\varepsilon}+1)$. ∎
The hockey-stick divergence is defined as in (7) for $\gamma\geq 1$. Now, similar to the classical definition of the hockey-stick divergence (see [46, Eq. (12)]), we extend its definition to $\gamma\geq 0$ as follows:
@@
where $\rho$ and $\sigma$ are states and $(x)_{+}\coloneqq\max\{0,x\}$. This definition recovers (7) because $(1-\gamma)_{+}=0$ for $\gamma\geq 1$.
@@
Furthermore, using the monotonicity of $\gamma^{\prime}\mapsto E_{\gamma^{\prime}}$ for $\gamma^{\prime}\geq 1$, we find the following for $\gamma\leq e^{-\varepsilon}$:
@@
Next, we quantify the contraction of the hockey-stick divergence for $\gamma\in[e^{-\varepsilon},1]$, similar to Theorem 1 for $\gamma\in[1,e^{\varepsilon}]$.
@@
Let $\gamma\in[e^{-\varepsilon},1]$, $\mathcal{A}\in\mathcal{B}^{\varepsilon}$, and $\rho$ and $\sigma$ states. Then, we have that
@@
where $E_{\gamma}(\cdot\|\cdot)$ is defined in (56). Furthermore, for $\gamma\in[e^{-\varepsilon},1]$, the privatized contraction coefficient of hockey-stick divergence, $\eta^{\varepsilon}_{E_{\gamma}}$, satisfies the following inequality:
@@
Since $\gamma\in[e^{-\varepsilon},1]$, it follows that $1/\gamma\in[1,e^{\varepsilon}]$. By applying Lemma 1, we obtain the following:
@@
Theorem 1 and Corollary 1 strengthen the previously known upper bound on the privatized contraction coefficient for the normalized trace distance (i.e., $\eta^{\varepsilon}_{T}$ by choosing $\gamma=1$), as recalled in (40). Next, Theorem 2 states that the upper bound from Theorem 1, for trace distance, is indeed tight. Under $\varepsilon$-classical local differential privacy (recall Definition 1), Theorem 2 was presented as [7, Corollary 2.9]. Our method of proof is different from that presented for [7, Corollary 2.9].
@@
$D_{f}(\rho\|\sigma)$ $=\int_{1}^{\infty}f^{\prime\prime}(\gamma)E_{\gamma}(\rho\|\sigma)\,\mathsf{d}\gamma+\int_{0}^{1}f^{\prime\prime}(\nu)E_{\nu}(\rho\|\sigma)\,\mathsf{d}\nu$ (107)
@@
Proposition 6.
For $\varepsilon\geq 0$, the following equality holds:
$\sup_{\begin{subarray}{c}A\in\mathcal{B}^{\varepsilon},\\ \rho,\sigma\in\mathcal{D}\end{subarray}}D_{f}\!\left(\mathcal{A}(\rho)\|\mathcal{A}(\sigma)\right)=\frac{f\!\left(e^{\varepsilon}\right)+e^{\varepsilon}f\!\left(e^{-\varepsilon}\right)}{e^{\varepsilon}+1},$ (113)
where $D_{f}(\cdot\|\cdot)$ is defined in (102) and the optimization is over all $\varepsilon$-QLDP channels $\mathcal{A}$ and states $\rho$ and $\sigma$. The equality is achieved by a pair of orthogonal states and a channel $\mathcal{A}$ with the output dimension two.
Proof:
See Section -C. ∎